# From chimeras to extensive chaos in networks of heterogeneous Kuramoto oscillator populations

### v1

@@ VI Erdős-Rényi networks
Figure 8: Approximated region of transverse instabilities of the homogeneous state in Erdős-Rényi networks with average degree $\overline{d}=10$. Black-shaded region indicates where at least one mode is unstable (i.e., $\text{Re}[\lambda_{2}^{(+)}]>0$). Purple-shaded region indicates where $N-1$ modes are unstable. Vertical dashed line indicates the Kuramoto synchronization transition H0.
@@
Here we discuss the dynamics that emerge in irregular network topologies. In order to simplify the analysis, we restrict ourselves on connectivities derived from Erdős-Rényi networks with fixed average degree $\overline{d}=10$. Let $A=(a_{\sigma\kappa})$ be the adjacency matrix of an (undirected) Erdős-Rényi network (i.e., $a_{\sigma\kappa}\in\{1,0\}$). Then, we consider a connectivity matrix $C=(c_{\sigma\kappa})$ given by
@@
$c_{\sigma\kappa}=\begin{cases}\mu\quad&\text{ if }\quad\sigma=\kappa\\ \displaystyle{\frac{1-\mu}{d_{\sigma}}}a_{\sigma\kappa}\quad&\text{ if }\quad\sigma\neq\kappa.\end{cases}$ (20)
@@
where $d_{\sigma}$ is the degree of node $\sigma$ in the Erdős-Rényi model. This connectivity preserves the row-normalization required for the existence of homogeneous states, as well as introduces the self-coupling strength parameter $\mu$.
@@
Unlike the ring scenario, we are unaware of explicit analytical results on the spectra of $C$ (which does not need to coincide with the spectra of the adjacency matrix $A$). However, combining analytical arguments with a numerical exploration, we can provide an accurate estimation of the region of transverse instabilities. By the Gershgoring circle theorem Wilkinson (1967), the eigenvalues of $C$ are contained in the interval $\Lambda_{k}\in[1,2\mu-1]$. Therefore, based on Fig. 1(b), transverse instabilities of the homogeneous state will occur. The shape of the region will depend on the second largest eigenvalue $\Lambda_{2}$. If $\tilde{\Lambda}$ is an eigenvalue of the matrix $D^{-1}A=(a_{\sigma\kappa}/d_{\sigma})$, then $\Lambda=\mu+(1-\mu)\tilde{\Lambda}$ is an eigenvalue of $C$. Averages over $100$ simulated networks with different sizes show that $\tilde{\Lambda}$ are densely distributed in the interval [-0.6,0.6]. In particular, $\tilde{\Lambda}_{2}\approx 0.6$ as $N$ increases, and therefore $\Lambda_{2}\approx 0.4\mu+0.6$. Using this approximation in the MSF Eq. (14) provides the region of transverse instabilities for ER networks in the $(K,\mu)$ parameter space. Black region in Fig. 8 shows the region where at least one mode of the homogeneous state is unstable. Also in this case the number of unstable directions in the region increases with $N$, with the purple region in Fig. 8 indicating the parameter values for which the homogeneous state has a maximum number of unstable modes ($N-1$).
@@
In order to capture the emergent dynamics, Figure 6(a3-d3) displays the results of simulations performed with fixed $\mu=0.9$ and different values of $K$ for an ER network of $N=128$ populations. As in the ring networks, the numerical outcome shows a scenario of consistent high-dimensional chaos for most of the transverse instability region (indicated by vertical dashed lines). Interestingly, two different types of complex dynamics seem to emerge: First, the level of heterogeneity captured by $\mathcal{S}$ shows a region of high dynamical variety among nodes (see red-shaded region in panel (a3), $6\lesssim K\lesssim 8$). In this case the Lyapunov spectra shows many exponents with similar positive values (panel (b3)), transducing to a very high Kaplan-Yorke dimension (panel (c3)) and dynamical entropy (panel (d4)). At $K\approx 8$, all measures show a significant reduction of complexity, indicating a transition to a different chaotic regime from $8\lesssim K\lesssim 20$. In this case, the patterns lose heterogeneity, but the Lyapunov spectra still shows a regime of high-dimensional chaos. In fact, the number and value of the largest exponents increase monotonically with $K$, leading to an equally gradual increase of $\mathcal{D}_{\text{KY}}$ and $h_{\text{KS}}$ up to the stabilization of the homogeneous state ($K\approx 20$).
@@
As in the ring model, we test the robustness of the chaotic dynamics upon increasing network size. Green dots in Figure 7(a) show the entire Lyapunov spectra for $K=15$ and $\mu=0.9$ upon increasing system size. Apart from some deviations for the smaller network ($N=64$), the different spectra show a good collapse. Moreover, green symbols in Fig.7(b) and (c) indicate that both fractal dimension and dynamical entropy increase linearly with $N$, confirming thus the extensive nature of the chaotic regime in ER topologies.
@@ VII Conclusions
Recent works have studied the dynamics of networks of identical Kuramoto-Sakaguchi populations by means of the Ott-Antonsen equations under differerent parameter setups, coupling schemes, and system sizesAbrams et al. (2008); Laing (2009a, b, 2012); Martens (2010a, b); Lee and Krischer (2023b, a); Bick et al. (2020). This paper provides a leap forward in these efforts by providing a general expression for the instabilities of the homogeneous state as a function of the connectivity eigenmodes. We proved that, in networks without disorder ($\Delta=0$), the homogeneous state is always transversally stable regardless the topology. On the other hand, for disordered populations ($\Delta>0$), transverse instabilities occur. These instabilities, which lead to chimera states in two-population networksLaing (2009a); Kotwal, Jiang, and Abrams (2017); Guo et al. (2021), give rise to extensive chaos in large nearest-neighbour ring and random connectivity schemes.
@@
The emergence of extensive chaos in ring networks could apparently conflict with the findings of Laing Laing (2023), which studied a similar setup but with non-local connectivity via a discrete kernel. In their work, the row-normalization of the connectivity matrix is preserved, thus our theoretical framework also applies. The eigenvalues of their connectivity matrix can be analytically computed, and we found they are all zero except for three: $\Lambda_{1}=1$, $\Lambda_{2}=\Lambda_{3}>0$. In particular, $\lim_{N\to\infty}\Lambda_{2}=B/2$ where $B=0.35$. This localized spectra differs largely from the spectra of nearest neighbour ring topologies given by Eq. (18), which covers the interval $[2\mu-1,1]$ densely for large $N$. Although a more in-detail comparison between both setups is out of the scope of this paper, it is likely that transverse instabilitites in [Laing, 2023] only occur through $\Lambda_{2}$ and $\Lambda_{3}$ rather than several concurrently unstable eigenmodes. This might explain the disappearance of chaos for large networks in that setup.
@@
These differences between non-local and local interactions in rings indicate that network topology and, specifically, their spectra, play an important role on the properties of the resulting dynamical states. Here, we restricted ourselves on simple structures, without an in-depth analyzis on how specific topological features affect the emergent behavior. Future work should analyze the effect of varying network density in both ring and Erdős-Rényi networks, or the impact of other topologies such as scale-free or small world.
@@
On a different note, the extensive nature of the chaotic states and the patterns displayed in ring networks are an indication that this form of space-time chaos could correspond to a turbulent state. Moreover, the dependence of the growth rates on the structural eigenmodes depicted in Fig. 1(a) resembles that of the Benjamin-Feir instability in the complex Ginzburg-Landau equationBenjamin and Feir (1967). Similar turbulent states have been studied in arrays of individual Kuramoto-Sakaguchi oscillators with non-local coupling Wolfrum, Gurevich, and Omel’chenko (2016); Bolotov et al. (2018, 2021). Further analysis should thus address to what extent the properties of our setup for increasing $N$ match those of phase or amplitude turbulence in spatially extended systems Kuramoto (1984).
@@
Here we have studied the dynamics of the Kuramoto-Sakaguchi populations at the thermodynamic limit via the Ott-Antonsen equations. The global attractiveness of this low-dimensional manifold for $\Delta>0$ has been recently provenCestnik and Pikovsky (2022), strengthening the validity of this approach to study networks of Kuramoto-Sakaguchi populations. Nonetheless, previous works have shown that (low-dimensional) chaotic dynamics unveiled by the Ott-Antonsen equations might display different properties in finite-size simulations Laing (2012). A detailed comparison of simulations of the microscopic system (1) with finite $M$, and those of the Ott-Antonsen description Eq. (3), could provide valuable insights to understanding the convergence of Kuramoto-Sakaguchi ensembles to the low-dimensional manifold. Nonetheless, such study would require a large computational effort due to the large number of degrees of freedom involved.
@@
Finally, although our study does not stem from any particular physical setup, the Kuramoto-Sakaguchi model successfully describes a wide variety of real-world phenomena, including power gridsFilatrella, Nielsen, and Pedersen (2008), circadian rhythmsStrogatz (1987), and biofilmsLiu et al. (2017). A particularly explored application is neural dynamics, where the Kuramoto-Sakaguchi model has been rigorously derived from spiking neuron modelsClusella, Pietras, and Montbrió (2022). In this context, the mean-field description provided by the Ott-Antonsen equations offers a simplified neural mass modelBick et al. (2020). Thus, our analysis of the coupled system Eq. (3) provides a simple, yet grounded, framework to analyze whole brain dynamics where each brain region is represented by a Kuramoto-Sakaguchi population. In fact, the study of transverse instabilities in whole brain models is a powerful tool to unveil the multiscale dynamics of these systems, even when the single-node dynamics evolve according to generic neural mass modelsClusella et al. (2023).
@@ VIII Acknowledgements
The authors would like to thank Diego Pazó for useful feedback on this work.
@@ Appendix B Numerical simulations and numerical continuation
The code used in this paper is openly available at github.com/pclus/kuramoto-sakaguchi-network. Simulations of the model and computation of Lyapunov exponents have been performed in Julia via the DynamicalSystems.jl package Datseris (2018). In particular we have employed the Runge-Kutta 4 algorithm with a fixed time step of $dt=10^{-2}$.
Simulations of the system employed an initial transient evolution of $6000$ time units. This is a conservative transient time set to achieve a good convergence near the bifurcations, but for most parameter values shorter transients provide the same results. After this transient, we simulate both, phase space and tangent space dynamics for the computation of the Lyapunov exponents. The tangent space is initially evolved for 200 time units, after which we start computing all the Lyapunov exponents for a total of 1000 time units. The QR decomposition for computation of the spectra is invoked at every time unit (i.e., $1/dt$ time steps).
Simulations of the the ring network have been initialized at the antiphase state with an additional small random perturbation. This prevents the system from stalling into a long chaotic transient in the region where the antiphase state is stable. For the random network, since no trivial antiphase state exists, simulations have been initialized at the homogeneous state with a random small perturbation.
Bifurcation diagrams of Fig. 2 have been partially obtained with auto-07p numerical continuation softwareDoedel et al. (2007). In order to simplify the analysis, the 2-population model with the connectivity matrix (16), it has been rewritten as a three-dimensional system in real space by considering the polar representation of the Kuramoto complex parameters $Z_{1}=R_{1}e^{i\Phi_{1}}$ and $Z_{2}=R_{2}e^{i\Phi_{2}}$ and defining a new variable $\Theta=\Phi_{1}-\Phi_{2}$. Then, Eq. (3) reads
@@
The auto-07p codes for the numerical continuation and a tutorial guide to obtain some of the bifurcations is openly available at github.com/pclus/kuramoto-sakaguchi-network.
@@ Appendix C Parameter reduction
System (3) contains three parameters: $\Delta$, $K$, and $\alpha$. It is possible to rescale the system in order to reduce the parameter number to two. Some worksLaing (2009a); Kotwal, Jiang, and Abrams (2017); Guo et al. (2021) consider systems in which time is rescaled as $\tilde{t}=Kt$ so that the coupling strength $K$ can be removed from the equations or, equivalently, let it constant as $K=1$. In this case Eq. (3) reads
$\frac{{\rm d}Z_{\sigma}}{{\rm d}\tilde{t}}=-\tilde{\Delta}Z_{\sigma}+\frac{1}{2}\sum_{\kappa=1}^{N}c_{\sigma\kappa}\left(Z_{\kappa}e^{-i\alpha}-Z^{2}_{\sigma}Z^{*}_{\kappa}e^{i\alpha}\right)$ (26)
where $\tilde{\Delta}=\Delta/K$ and $K>0$.
@@
In this work we choose to not rescale any parameter in the system, so the analysis remains valid in all the extreme cases ($\Delta=0$ and $K=0$). Nonetheless, we are mostly interested in the heterogeneous case ($\Delta>0$), thus we let $\Delta=1$ without loss of generality. Therefore, in the 2-population model, the equivalence between the formulation of previous studies Laing (2009a, 2012); Kotwal, Jiang, and Abrams (2017); Guo et al. (2021) and that given in Eq. (3) with $\Delta=1$ is $D=1/K$, $\mu=p$, $\nu=1-p$, and $A=\mu-\nu=2p-1$.

### v2

@@ VI Erdős-Rényi networks
Figure 8: Approximated region of transverse instabilities of the homogeneous state in Erdős-Rényi networks with average degree $\overline{d}=10$. The grey-shaded region indicates where at least one mode is unstable (i.e., $\text{Re}[\lambda_{2}^{(+)}]>0$). The purple-shaded region indicates where $N-1$ modes are unstable. The vertical dashed line indicates the Kuramoto synchronization transition H0.
@@
Here we discuss the dynamics that emerge in irregular network topologies. In order to simplify the analysis, we restrict ourselves on connectivities derived from Erdős-Rényi (ER) networks with fixed average degree $\overline{d}=10$. Let $A=(a_{\sigma\kappa})$ be the adjacency matrix of an (undirected) Erdős-Rényi network (i.e., $a_{\sigma\kappa}\in\{1,0\}$). Then, we consider a connectivity matrix $C=(c_{\sigma\kappa})$ given by
@@
$c_{\sigma\kappa}=\begin{cases}\mu\quad&\text{ if }\quad\sigma=\kappa\\ \displaystyle{\frac{1-\mu}{d_{\sigma}}}a_{\sigma\kappa}\quad&\text{ if }\quad\sigma\neq\kappa,\end{cases}$ (20)
@@
where $d_{\sigma}$ is the degree of node $\sigma$ in the Erdős-Rényi model. This connectivity preserves row-normalization, as required for the existence of homogeneous states, as well as introduces the self-coupling strength parameter $\mu$.
@@
Unlike the ring scenario, we are unaware of explicit analytical results on the spectra of $C$ (which does not need to coincide with the spectra of the adjacency matrix $A$). However, combining analytical arguments with a numerical exploration, we can provide an accurate estimation of the region of transverse instabilities. By the Gershgoring circle theorem Wilkinson (1967), the eigenvalues of $C$ are contained in the interval $\Lambda_{k}\in[1,2\mu-1]$. Therefore, based on Fig. 1(b), transverse instabilities of the homogeneous state will occur. The shape of the region will depend on the second largest eigenvalue $\Lambda_{2}$. If $\tilde{\Lambda}$ is an eigenvalue of the matrix $D^{-1}A=(a_{\sigma\kappa}/d_{\sigma})$, then $\Lambda=\mu+(1-\mu)\tilde{\Lambda}$ is an eigenvalue of $C$. Averages over $100$ simulated networks with different sizes (not shown) indicate that $\tilde{\Lambda}$ are densely distributed in the interval [-0.6, 0.6]. In particular, $\tilde{\Lambda}_{2}\approx 0.6$ as $N$ increases, and therefore $\Lambda_{2}\approx 0.4\mu+0.6$. Using this approximation in the MSF, Eq. (14), provides the region of transverse instabilities for ER networks in the $(K,\mu)$ parameter space. The grey-shadowed region in Fig. 8 shows the region where at least one mode of the homogeneous state is unstable. Also in this case, the number of unstable directions in the region increases with $N$, with the purple region in Fig. 8 indicating the parameter values for which the homogeneous state has a maximum number of unstable modes ($N-1$).
@@
In order to capture the emergent dynamics, Figure 6(a3-d3) displays the results of simulations performed with fixed $\mu=0.9$ and different values of $K$ for an ER network of $N=128$ populations. As in the ring networks, the numerical outcome shows a scenario of consistent high-dimensional chaos for most of the transverse instability region (indicated by vertical dashed lines). Interestingly, two different types of complex dynamics seem to emerge: First, the level of heterogeneity captured by $\mathcal{S}$ shows a region of high dynamical variety among nodes (see red-shaded region in Fig. 6(a3), $6\lesssim K\lesssim 8$). In this case the Lyapunov spectra shows many exponents with similar positive values (panel b3), which translates into a very high Kaplan-Yorke dimension (panel c3) and dynamical entropy (panel d4). At $K\approx 8$, all measures show a significant reduction of complexity, indicating a transition to a different chaotic regime from $8\lesssim K\lesssim 20$. In this case, the patterns lose heterogeneity, but the Lyapunov spectra still shows a regime of high-dimensional chaos. In fact, the number and value of the largest exponents increase monotonically with $K$, leading to an equally gradual increase of $\mathcal{D}_{\text{KY}}$ and $h_{\text{KS}}$ until the homogeneous state becomes stable again (at $K\approx 20$).
@@
As in the ring model, we now test the robustness of the chaotic dynamics upon increasing network size. The green dots in Figure 7(a) show the entire Lyapunov spectra for $K=15$ and $\mu=0.9$ upon increasing system size. Overall, the different spectra collapse to a single well-defined curve. Moreover, the green symbols in Fig. 7(b,c) indicate that both the fractal dimension and the dynamical entropy increase linearly with $N$, confirming thus the extensive nature of the chaotic regime in ER topologies.
@@ VII Conclusions
Recent works have studied the dynamics of networks of identical Kuramoto-Sakaguchi populations by means of the Ott-Antonsen equations under differerent parameter setups, coupling schemes, and system sizesAbrams et al. (2008); Laing (2009a, b, 2012); Martens (2010a, b); Lee and Krischer (2023b, a); Bick et al. (2020). This paper provides a leap forward in these efforts by providing a general expression for the instabilities of the homogeneous state as a function of the eigenmodes of the structural connectivity matrix. We proved that, in networks without disorder ($\Delta=0$), the homogeneous state is always transversally stable regardless the topology. On the other hand, for disordered populations ($\Delta>0$), transverse instabilities occur. These instabilities, which lead to chimera states in two-population networksLaing (2009a); Kotwal, Jiang, and Abrams (2017); Guo et al. (2021), give rise to extensive chaos in large nearest-neighbour ring and random connectivity schemes.
@@
The emergence of extensive chaos in ring networks could apparently conflict with the findings of Laing Laing (2023), which studied a similar setup but with non-local connectivity via a discrete kernel. In their work, the row-normalization of the connectivity matrix is preserved, thus our theoretical framework also applies. The eigenvalues of their connectivity matrix can be analytically computed, and we found they are all zero except for three: $\Lambda_{1}=1$, $\Lambda_{2}=\Lambda_{3}>0$. In particular, $\lim_{N\to\infty}\Lambda_{2}=B/2$ where $B=0.35$. This localized spectra differs largely from the spectra of nearest neighbour ring topologies given by Eq. (18), which covers the interval $[2\mu-1,1]$ densely for large $N$. Although a more in-detail comparison between both setups is out of the scope of this paper, it is likely that the transverse instabilitites reported in Ref. [Laing, 2023] only occur through $\Lambda_{2}$ and $\Lambda_{3}$, rather through than several concurrently unstable eigenmodes. This might explain the disappearance of chaos for large networks in that setup.
@@
These differences between non-local and local interactions in rings indicate that network topology and, specifically, their spectra, play an important role on the properties of the resulting dynamical states. Here we restricted ourselves on simple structures, without an in-depth analysis of how specific topological features affect the emergent behavior. Future work should analyze the effect of varying network density in both ring and Erdős-Rényi networks, and the impact of other topologies such as scale-free or small world.
@@
On a different note, the extensive nature of the chaotic states and the patterns displayed in ring networks are an indication that this form of space-time chaos could correspond to a turbulent state. Interestingly, the dependence of the growth rates on the structural eigenmodes depicted in Fig. 1(a) resembles that of the Benjamin-Feir instability in the complex Ginzburg-Landau equationBenjamin and Feir (1967). Similar turbulent states have been studied in arrays of individual Kuramoto-Sakaguchi oscillators with non-local coupling Wolfrum, Gurevich, and Omel’chenko (2016); Bolotov et al. (2018, 2021). Further analysis should thus address to what extent the properties of our setup for increasing $N$ match those of phase or amplitude turbulence in spatially extended systems Kuramoto (1984).
@@
Here we have studied the dynamics of the Kuramoto-Sakaguchi populations at the thermodynamic limit via the Ott-Antonsen equations. The global attractiveness of this low-dimensional manifold for $\Delta>0$ has been recently provenCestnik and Pikovsky (2022), strengthening the validity of this approach to study networks of Kuramoto-Sakaguchi populations. Nonetheless, previous works have shown that (low-dimensional) chaotic dynamics unveiled by the Ott-Antonsen equations might display different properties in finite-size simulations Laing (2012). A detailed comparison of simulations of the microscopic system (1) with finite $M$, and those of the Ott-Antonsen description Eq. (3), could provide valuable insights to understanding the convergence of Kuramoto-Sakaguchi ensembles to the low-dimensional manifold. Nonetheless, such study would require a large computational effort, due to the large number of degrees of freedom involved.
@@
Finally, although our study does not stem from any particular physical setup, the Kuramoto-Sakaguchi model successfully describes a wide variety of real-world phenomena, including power gridsFilatrella, Nielsen, and Pedersen (2008), circadian rhythmsStrogatz (1987), and bacterial biofilmsLiu et al. (2017). A particularly explored application is neural dynamics, where the Kuramoto-Sakaguchi model has been rigorously derived from spiking neuron modelsClusella, Pietras, and Montbrió (2022). In this context, the mean-field description provided by the Ott-Antonsen equations leads to a simplified neural mass modelBick et al. (2020). Thus, our analysis of the coupled system Eq. (3) provides a simple, yet grounded, framework to analyze whole brain dynamics where each brain region is represented by a Kuramoto-Sakaguchi population. In fact, the study of transverse instabilities in whole brain models is a powerful tool to unveil the multiscale dynamics of these systems, even when the single-node dynamics evolve according to generic neural mass modelsClusella et al. (2023).
@@ VIII Acknowledgements
The authors would like to thank Diego Pazó and Alessandro Torcini for useful feedback on this work. PC and JGO have received funding from the Future and Emerging Technologies Programme (FET) of the European Union’s Horizon 2020 research and innovation programme (project NEUROTWIN, grant agreement No 101017716). JGO was also financially supported by the Spanish Ministry of Science and Innovation, the Spanish State Research Agency and FEDER (project reference PID2021-127311NB-I00), and by the ICREA Academia program. This work has done while PF was a student at the ETSETB, Universitat Politècnica de Catalunya.
@@ Appendix B Numerical simulations and numerical continuation
The code used in this paper is openly available at [Clusella, 2024] (also accessible through https://github.com/pclus/KuramotoPopulationNetwork). Simulations of the model and computation of Lyapunov exponents have been performed in Julia via the DynamicalSystems.jl package Datseris (2018). In particular we have employed the Runge-Kutta 4 algorithm with a fixed time step of $dt=10^{-2}$.
Simulations of the system employed an initial transient evolution of $6000$ time units. This is a conservative transient time set to achieve a good convergence near the bifurcations, but for most parameter values shorter transients provide the same results. After this transient, we simulate both phase space and tangent space dynamics for the computation of the Lyapunov exponents. The tangent space is initially evolved for 200 time units, after which we start computing all the Lyapunov exponents for a total of 1000 time units. The QR decomposition for computation of the spectra is invoked at every time unit (i.e. $1/dt$ time steps). The computation of the Kaplan-Yorke formula requires a tolerance to discern zero and non-zero exponents, which we set to $10^{-4}$.
The results depicted in Fig. 6(a2-d2) require longer computation times due to the existence of the slow chaotic regime. In that case, the integration time step was chosen to be $dt=10^{-1}$, the transient in tangent space consisted of $10^{4}$ time units, and the Lyapunov exponents were computed for $3\cdot 10^{4}$ time units with the QR decomposition at every 10 time units. Also in this case, we lowered the tolerance for the Kaplan-Yorke formula to $2\cdot 10^{-5}$.
Simulations of the ring network for $\mu=0.9$ were initialized at the antiphase state with an additional small random perturbation. This prevents the system from stalling into a long chaotic transient in the region where the antiphase state is stable. For all other cases, since the antiphase state does not exist or is unstable, simulations were initialized at the homogeneous state with a random small perturbation.
The bifurcation diagrams of Fig. 2 were partially obtained with the numerical continuation software auto-07pDoedel et al. (2007). In order to simplify the analysis, the 2-population model with the connectivity matrix (16) was rewritten as a three-dimensional system in real space by considering the polar representation of the Kuramoto complex parameters $Z_{1}=R_{1}e^{i\Phi_{1}}$ and $Z_{2}=R_{2}e^{i\Phi_{2}}$, and defining a new variable $\Theta=\Phi_{1}-\Phi_{2}$. Then, Eq. (3) reads
@@
A tutorial guide to obtain some of the bifurcations using the numerical continuation capabilities of auto-07p is openly available in the published repositoryClusella (2024).
@@ Appendix C Parameter reduction
System (3) contains three parameters: $\Delta$, $K$, and $\alpha$. It is possible to rescale this system in order to reduce the parameter number to two. Some worksLaing (2009a); Kotwal, Jiang, and Abrams (2017); Guo et al. (2021) consider systems in which time is rescaled as $\tilde{t}=Kt$, so that the coupling strength $K$ can be removed from the equations or, equivalently, let it constant as $K=1$. In this case Eq. (3) reads
$\frac{{\rm d}Z_{\sigma}}{{\rm d}\tilde{t}}=-DZ_{\sigma}+\frac{1}{2}\sum_{\kappa=1}^{N}c_{\sigma\kappa}\left(Z_{\kappa}e^{-i\alpha}-Z^{2}_{\sigma}Z^{*}_{\kappa}e^{i\alpha}\right)$ (26)
where $D=\Delta/K$ and $K>0$.
@@
In this work we choose not to rescale any parameter in the system, so the analysis remains valid in the limit cases $\Delta=0$ and $K=0$. Nonetheless, we are mostly interested in the heterogeneous case ($\Delta>0$), thus we let $\Delta=1$ without loss of generality. Therefore, in the 2-population model, the equivalence between the formulation of previous studies Laing (2009a, 2012); Kotwal, Jiang, and Abrams (2017); Guo et al. (2021) and that given in Eq. (3) with $\Delta=1$ is simply $D=1/K$.